# Approximately-symmetric neural networks for quantum spin liquids

### v1

@@ III.2 Invariance error
Furthermore, we found that the invariance error $\tilde{\epsilon}$ significantly increases with the increasing values of the $h_{z}$ field. To investigate this behaviour further we plot the $\log\tilde{\epsilon}$ as a function of $h_{z}$ field in the Fig. S5a. First, as expected, at $h_{z}=0$ the approximately-symmetric architecture learns to stay fully invariant (despite of extra flexibility). Second, we observe that $\mathrm{d}\log\tilde{\epsilon}/\mathrm{d}h_{z}$ reveals a peak matching the critical point found by more conventional observables (as in Fig. 3 main text). Such behavior is most likely comes from the connection of $\epsilon$ invariance error to the extent by which $X$-Wilson loops act non-trivially on a system’s ground state: $\epsilon\sim||A_{v}|\psi\rangle-|\psi\rangle||$. It should be noted that, as expected, $\tilde{\epsilon}$ exhibits no such divergent behaviour when tuning symmetry-preserving terms (e.g., $h_{x}$ magnetic field) – under which the network stays fully invariant.

### v2

@@ Abstract
[…]
Appendix S4 Approximately-symmetric NQS for Rydberg spin liquids
Here we provide more information about approximately-symmetric neural network applied to a PXP limit of the Rydberg Hamiltonian on a ruby lattice. First, we efficiently incorporate the Rydberg blockade within each triangle by reducing the local Hilbert space to that of a spin-$3/2$, since double-occupancy configurations cost infinite energy in the PXP limit. The resulting local basis over a triangle is $\{|ggg\rangle,|egg\rangle,|geg\rangle,|gge\rangle\}$, effectively defining a spin-$3/2$ system on a honeycomb lattice. Second, we use an approximately-symmetric neural network consisting of a gauge invariant block, non-gauge-symmetric CNN added together with the non-gauge-symmetric mean-field ansatz (thus forming the “RPP” architecture discussed before). Relevant gauge invariant Wilson operators – with respect to which gauge-invariant product non-linearities are constructed – are shown in Fig. S7(a) and Fig. S7(b) - see also Ref. [4]. The mean-field ansatz $\textrm{MF}=\sum_{i}c_{i}n_{i}$ is added for numerical stability. We symmetrize the input to the neural network with lattice symmetries (horizontal and vertical parity). We use, single layer, 4204-20 features in the non-symmetric part (in addition to the mean-field features), GELU activation functions and stochastic reconfiguration with the SVD-based solution to the set of linear equations. Finally, on small system sizes we benchmark our architecture against RBMs enhanced with spin-$3/2$ blockade constraints and a mean-field ansatz respecting all lattice symmetries (including inversion around the $x$ and $y$ axes).
We find evidence of the quantum spin liquid phase by inspecting Wilson loop operators for large system sizes ($L\geq 6$) and finite size versions of BFFM order parameters (on $L=6$ they evaluate to $\langle X_{BFFM}\rangle\sim 0.01$ for equal-superposition probing loop passing through $12$ spins-$1/2$ and $\langle Z_{BFFM}\rangle\sim 0.07$ for Gauss’ law probing loop passing through $18$ spins-$1/2$), in agreement with the predictions from iDMRG [4]. Loops used to calculate BFFM order parameters are defined in Fig. S7(c).
Figure S7: (a) $X_{loop}=\prod_{j\in\textrm{hexagon}}\tilde{X}_{j}$ operators used to construct gauge invariant non-linearites for the ruby model. Red filled/empty circles denote Rydberg atoms in an excited/ground state (b) Action of $\tilde{X}_{j}$ within each triangle (c) Loop operators used to calculate BFFM order parameters on $L=6$.